Two stations A and B are simultaneously trying to access a satellite channel using random access protocol. A transmits a frame with probability p and B transmits a frame with probability 2p.

Derive the expression for channel utilization as a function of p and explain how that derivation was found.
Answer #1

Channel will be utilized if and only if exactly one of the station transmits a frame while other remains silent.

Hence channel utilization =( Probability that A transmits a frame)*(Probability that B does not transmit a frame) + (Probability that A does not transmit a frame)*(Probability that B transmit a frame) = p*(1-2p) + (1-p)*2p = p(1-2p + 2 - 2p) = p*(3-4p)

Please comment for any clarification